Suppose that A and B are events defined on a common sample space and that the following probabilities are known. Find P(A or B). (Give your answer correct to two decimal places.)
P(A) = 0.35, P(B) = 0.44, P(A | B) = 0.22

Answer by stanbon(75887)

Suppose that A and B are events defined on a common sample space and that the following probabilities are known. Find P(A or B). (Give your answer correct to two decimal places.)
P(A) = 0.35, P(B) = 0.44, P(A | B) = 0.22
-----
P(A and B) = P(A | B)*P(B) = 0.22*0.44 = 0.0968
------
Then, P(A or B) = P(A) + P(B) - P(A and B)
----
= 0.35+0.44-0.0968
----
= 0.693
------------
Cheers,
Stan H.
-------------